Samsung Aviation is considering purchasing equipment to manufacture a part of the wing fixture for an aircraft. The cost of the system is $5 million with life expectancy of 10 years and annual operating cost of $200,000. Samsung is using a MARR of 12 percent, expecting an annual revenue of $1.5 million for 10 years, and a terminal cash flow of $25,000. Using present worth analysis, is this investment economically justified?
1
Expert's answer
2020-05-04T11:58:05-0400

To evaluate if this investment is economically justified we should calculate NPV of this project:

"NPV = -5,000,000 + \frac{1,5000,000 - 200,000}{(1 + 0.12)^1} + \frac{1,5000,000 - 200,000}{(1 + 0.12)^2} + ... + \frac{1,5000,000 - 200,000 + 25,000}{(1 + 0.12)^{10}} = 2,353,339."

So, as NPV > 0, then this investment is economically justified.